OpenStudy (yanasidlinskiy):

STATS HELP PLEASE? A standard deck of cards contains 52 cards. One card is selected from the deck. ​(a) Compute the probability of randomly selecting a six or three. ​(b) Compute the probability of randomly selecting a six or three or nine. ​(c) Compute the probability of randomly selecting a queen or diamond.

OpenStudy (yanasidlinskiy):

There's 4 6's and 4 3's What's my next step in solving this problem?

OpenStudy (yanasidlinskiy):

Would I divide 6/52*3/52?

OpenStudy (jadzia):

I think the problem is composed of mutually exclusive events, so if I were you I would add individual probabilities. Not 100% sure

OpenStudy (anonymous):

a: 8/52 b: 12/52 c: 13+3=16, 16/52

OpenStudy (yanasidlinskiy):

Can you explain that?

OpenStudy (jadzia):

Probability of getting a six: 4/52 Probability of getting a three: 4/52 P(6 or 3) = 4/52 + 4/52 = 8/52 or 2/13 Same thing you have to do with the rest.

OpenStudy (yanasidlinskiy):

Thanks:)

OpenStudy (anonymous):

Sorry, this site is being unstable and slow to respond. Do not know if that is intended by the management or it's an act of God.

OpenStudy (yanasidlinskiy):

LOL. Maybe an act of God? hahha